Any line parallel to 4y-x=8 must have a slope equal to______________.
Answer by nerdybill(7384)

Any line parallel to 4y-x=8 must have a slope equal to______________.
Put
4y-x=8
in "slope-intercept" form:
y = mx + b
.
4y-x=8
4y = x+8
y = (1/4)x + 2
slope is 1/4
.
Any line parallel to 4y-x=8 must have a slope equal to 1/4